solve the solution by substitution:
y=11-2x
x=18-3y

Answer by mananth(16949)

y=11-2x
x=18-3y
plug value of y n the second equation
x=18-3(11-2x)
x=18-33+6x
x=-15+6x
add -6x
-6x+x=-15
-5x=-15
x=-15/-5
x=3
....
plug value of x in the first equation
y=11-2*3
y=11-6
y=5
....